# Identification by non-Gaussianity in structural threshold and smooth transition vector autoregressive models

### v4

@@ B.4 Proof of Proposition 2
Note that by Assumption A.1(II), Assumption 1 in Saikkonen (2008) is satisfied. Moreover, by the constraint $\sum_{m=1}^{M}\alpha_{m,t}=1$, Assumption 2(a) in Saikkonen (2008) satisfied, and as we assume that the set $\text{l}_{2}$ in Saikkonen (2008) is an empty set, also Assumption 2(b) is satisfied. Consequently, Condition (10) in Saikkonen (2008) is satisfied and by our Assumption A.1(I) also Condition (19). The difference of our setup to the conditions of Theorem 1 of Saikkonen (2008) is that our STVAR model specifies the conditional covariance matrix differently to the STVEC model defined in Equation (8) of Saikkonen (2008).
@@
That is, the first two terms on the right side of Equation (17) in Saikkonen (2008), reduce to the first term on the right hand side of Equation (3.2) defining our structural STVAR model. The third term on the right side of Equation (17) in Saikkonen (2008), however, reduces to $(\sum_{m=1}^{M}\alpha_{m,t}\Omega_{m})^{1/2}e_{t}$, where $\Omega_{1},...,\Omega_{M}$ are positive definite $(d\times d)$ covariance matrices of the regimes. This is different to the second term of Equation (3.2), which takes the form $\sum_{m=1}^{M}\alpha_{m,t}B_{m}e_{t}$ when the impact matrix (3.3) is substituted to it. Nonetheless, it turns out that with only slight modifications, the proof of Theorem 1 in Saikkonen (2008) applies to our model as well.
@@
The conditional covariance matrix is relevant in only a few places in the proof of Theorem 1 in Saikkonen (2008). Firstly, on page 312, Saikkonen (2008) makes use of the positive definiteness of the conditional covariance matrix $H(S^{\prime}JZ_{t-1})$ (or $\sum_{m=1}^{M}\alpha_{m,t}\Omega_{m}$ in our notation) to conclude that the smallest eigenvalue of $H(S^{\prime}Jx)$, $x\in\mathbb{R}^{dp}$, is bounded away from zero in compact subsets of $\mathbb{R}^{dp}$. Saikkonen (2008) argues that thereby by Theorem 2.2(ii) of Cline and Phu (1998), the Markov chain $Z_{t}$ is an irreducible and aperiodic T-chain on $\mathbb{R}^{dp}$, which is then used to conclude further results. Concerning our model, note that by Lemma 1, $B_{y,t}$ is invertible for all $t$ almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$, implying that the conditional covariance matrix $B_{y,t}(\boldsymbol{y}_{t-1})B_{y,t}(\boldsymbol{y}_{t-1})^{\prime}$ is positive definite, where the dependence of the conditional covariance matrix on the $(dp\times 1)$ random vector $\boldsymbol{y}_{t-1}=(y_{t-1},...,y_{t-p})$ is made explicit in the notation. Hence, the smallest eigenvalue of $B_{y,t}(x)B_{y,t}(x)^{\prime}$, $x\in\mathbb{R}^{dp}$, is bounded away from zero in compact subsets of $\mathbb{R}^{dp}$. It follows from this, Equation (3.2), and the invertibility of $B_{y,t}$ that we can apply Theorem 2.2(ii) of Cline and Phu (1998) to conclude that the Markov chain $\boldsymbol{y}_{t}$ is an irreducible and aperiodic T-chain on $\mathbb{R}^{dp}$, similarly to the process $Z_{t-1}$ in Saikkonen (2008).
Secondly, the conditional covariance matrix appears in the proof of Saikkonen (2008) on page 313, where Saikkonen (2008) shows that an appropriate version of Condition (15.3) of Meyn and Tweedie (1993, p. 358) is satisfied. To show that this condition holds for our model, we need to show that an appropriate version of Equation (A.5) in Saikkonen (2008) holds for our model, i.e., that the term $E[\sup_{A\in\mathcal{A}^{j}}||AB_{y,t}(x)e_{t}||^{2}]$ is bounded from above by a finite real number, where $j$ is a positive integer, $\mathcal{A}=\{\boldsymbol{A}_{1},...,\boldsymbol{A}_{M}\}$, and $\boldsymbol{A}_{1},...,\boldsymbol{A}_{M}$ as well as the set $\mathcal{A}^{j}$ are defined in Appendix A.
@@
$\sup\{||A||:A\in\cup_{j=1}^{N}\mathcal{A}^{j}\}<c,$ (B.30)
where |||||$\cdot$||| | ⋅ | | is Eucledian norm, $N$ is a positive integer, and $c$ a finite real number. Hence, because $B_{y,t}(x)$ is a constant here, by Jensen’s inequality and submultiplicativity of norms, we have
$E\left[\sup_{A\in\mathcal{A}^{j}}||AB_{y,t}(x)e_{t}||^{2}\right]\leq\sup_{A\in\mathcal{A}^{j}}||A||^{2}||B_{y,t}(x)||^{2}E[||e_{t}||]^{2},$ (B.31)
where the right side is bounded from above by some finite constant (by (B.30) and the finite second moment of $e_{t}$ given by Assumption 1). Thus, an appropriate version of Equation (A.5) in Saikkonen (2008) holds for our model. Since the rest of the proof of Theorem 1 in Saikkonen (2008) thereby applies to our model as well, this concludes the proof.∎
Appendix C Penalized nonlinear least squares estimation of STVAR models
@@
C.1 Nonlinear least squares estimation
@@
$Q(\alpha)=\sum_{t=1}^{T}u_{t}(\alpha)^{\prime}u_{t}(\alpha),$ (C.1)
@@
where $u_{t}(\alpha)=y_{t}-\sum_{m=1}^{M}\alpha_{m,t}\mu_{m,t}$ and the parameter $\alpha$ contains the weight function parameters. [^5] With threshold weights (2.3), $\alpha=(r_{1},...,r_{M-1})$, and with logistic weights (2.4) weights, $\alpha=(c,\gamma)$. If the transition weights are exogenous, we simply treat them as known constants. The NLS estimates of the parameters $\phi_{m},A_{m,j}$, $j=1,...,p$, corresponding to the weight function parameter $\alpha$ are obtained by minimizing $Q(\alpha)$ with respect to these parameters.
@@
If the transition weights $\alpha_{1,t},...,\alpha_{M,t}$ are not exogenous, also the parameter $\alpha$ (they depend on) needs to be estimated. Since the NLS estimates of the AR parameters are analytically obtained for given $\alpha$, the optimization can be performed with respect to $\alpha$ only. However, this optimization problem is multimodal, and thereby standard gradient based optimization methods that are not able escape from local maxima are not sufficient. Since we only need approximate estimates to serve as starting values for Step 2 and 3 estimation in the three-step procedure described in Appendix 4.3, the following procedure is adopted.
@@
First, determine the values of the transition weights $\alpha_{m,t}$, $m=1,...,M$, $t=1,...,T$, that can be deemed reasonable for estimation, i.e., transition weights that allocate a sufficient enough contribution to each regime in the model. As an extreme example, transition weights with $\alpha_{1,t}=0$ for all $t=1,...,T$ are not reasonable, as Regime 1 would not be involved in the model at all. Thus, each regime should contribute to the model to an extend that is reasonable enough for estimation.
@@
One way to assess the contribution of each regime is by summing over the transition weights, i.e., calculating the sums $\dot{\alpha}_{m}\equiv\sum_{t=1}^{T}\alpha_{m,t}$, $m=1,...,M$. In the special case of discrete transition weights such as the threshold weights (2.3), the sums $\dot{\alpha}_{m}$ express the numbers of observations from each regime. Therefore, one way to determine whether a regime contributes enough to the model is to check whether the sums $\dot{\alpha}_{m}$ are large enough compared to the number of parameters in that regime for all $m=1,...,M$. Denote the large enough sum as $\tilde{T}$. We assume that if there are $k$ parameters in a regime, and there should be at least three times as many observations from all the $d$ variables combined, implying $\tilde{T}=3kd^{-1}$, then transition weights that satisfy $\dot{\alpha}_{m}\geq\tilde{T}$ for all $m=1,...,M$ are deemed reasonable enough. [^6] We acknowledge that our assumption of at least three times as many observations from all $d$ variables combined as there are parameters in a regime is somewhat arbitrary, but it attempts to deter obvious cases of overfitting, while still allowing flexibility when the model is large compared to the number of observations.
@@
Second, specify a range of reasonable parameter values for each scalar parameter in the $(s\times 1)$ vector $\alpha$, and specify a grid of parameter values within that range. For instance, for a logistic STVAR model, the location parameter should, at least, lie between the lowest and greatest value in the observed series of the switching variable, whereas the scale parameter can range from a small positive number some large positive number that implies almost discrete regime-switches. With $n$ grid points for each $s$ scalar parameters in $\alpha$, the number of different parameter vectors $\alpha$ with values in the grid points is $n^{s}$. Hence, we recommend selecting a sufficiently small number of grid points $n$, so that estimation can be performed in a reasonable computation time. Denote the $i$th vector of $\alpha$ containing parameter values in the grid points as $\alpha^{(i)}$, $i=1,...,n^{s}$. Then, obtain the subset of the grid points $\alpha^{(i)}$ that satisfy $\dot{\alpha}_{m}\geq\tilde{T}$ for all $m=1,...,M$. Denote this subset as $\mathcal{W}$.
@@
Finally, calculate the NLS estimates of the AR parameters for all $\alpha^{(i)}\in\mathcal{W}$, and the estimates yielding the smallest sum of squares of residuals $Q(\alpha)$ are the approximate NLS estimates of the parameters $\phi_{1},....,\phi_{M}$, $\varphi_{1},...,\varphi_{M}$, and $\alpha$. Notably, the NLS estimation does not necessarily produce estimates that satisfy the usual stability condition for all or any of the regimes.
C.2 Penalized nonlinear least squares estimation
@@
Penalized nonlinear least squares estimation of STVAR models works similarly to the NLS estimation, except that instead of the sum of squares function (C.1), the penalized sum of squares function is minimized. We define the penalized sum of squares function as
$PQ(\alpha)=Q(\alpha)-P(\alpha),$ (C.2)
@@
$P(\alpha)=\kappa\widehat{RSS}\sum_{m=1}^{M}\sum_{i=1}^{dp}\text{max}\{0,|\rho(\boldsymbol{A}_{m}(\boldsymbol{\theta}))_{i}|-(1-\eta)\}^{2},$ (C.3)
@@
$\widehat{RSS}=\underset{\alpha\in\mathcal{W}}{\min}Q(\alpha)$ (C.4)
@@
is the minimized residual sum of squares, $|\rho(\boldsymbol{A}_{m}(\boldsymbol{\theta}))_{i}|$ is the modulus of the $i$th eigenvalue of the companion form AR matrix of Regime $m$, the tuning parameter $\eta\in(0,1)$ determines how close to the boundary of the stability region the penalization starts, and $\kappa>0$ determines the strength of the penalization. Whenever the companion form AR matrix of a regime has eigenvalues greater than $1-\eta$ in modulus, the penalization term (4.4) is greater than zero, and it increases in the modulus of these eigenvalues. The penalty term (C.3) is multiplied by $\widehat{RSS}$ to standardize the strength of the penalization with respect to the fit of the model.
@@
The penalized NLS estimates of $\phi_{m},A_{m,j}$, $j=1,...,p$, and $\alpha$, are obtained by first minimizing $Q(\alpha)$ for all $\alpha\in\mathcal{W}$, then calculating the penalty term $P(\alpha)$ for all $(\alpha)\in\mathcal{W}$, and finally selecting the estimates that give the smallest minimized the penalized sum of squares (C.2). In our empirical analysis in Section 5, we use the tuning parameter values $\eta=0.05$ and $\kappa=0.2$, in line with the tuning parameter values used in the PML estimation (see Section 4.2).
Appendix D A Monte Carlo study

### v5

@@ C.3 Proof of Proposition 1
As explained in Appendix B, our STVAR model is not nested to the STVEC model of Saikkonen (2008), and hence, additional steps are needed to show the applicability of the results of Saikkonen (2008) to our structural STVAR model. Note that by Assumption B.1(II), Assumption 1 in Saikkonen (2008) is satisfied. Moreover, by the constraint $\sum_{m=1}^{M}\alpha_{m,t}=1$, Assumption 2(a) in Saikkonen (2008) satisfied, and as we assume that the set $\text{l}_{2}$ in Saikkonen (2008) is an empty set, also Assumption 2(b) is satisfied. Consequently, Condition (10) in Saikkonen (2008) is satisfied and by our Assumption B.1(I) also Condition (19). The difference of our setup to the conditions of Theorem 1 of Saikkonen (2008) is that our STVAR model specifies the conditional covariance matrix differently to the STVEC model defined in Equation (8) of Saikkonen (2008).
@@
That is, the first two terms on the right side of Equation (17) in Saikkonen (2008), simplify to the first term on the right hand side of Equation (3.2) defining our structural STVAR model. The third term on the right side of Equation (17) in Saikkonen (2008), however, simplifies to $(\sum_{m=1}^{M}\alpha_{m,t}\Omega_{m})^{1/2}e_{t}$, where $\Omega_{1},...,\Omega_{M}$ are positive definite $(d\times d)$ covariance matrices of the regimes. This is different to the second term of Equation (3.2), which takes the form $\sum_{m=1}^{M}\alpha_{m,t}B_{m}e_{t}$ when the impact matrix (3.3) is substituted to it. Nonetheless, it turns out that with only slight modifications, the proof of Theorem 1 in Saikkonen (2008) applies to our model as well.
@@
The conditional covariance matrix is relevant in only a few places in the proof of Theorem 1 in Saikkonen (2008). Firstly, on Page 312, Saikkonen (2008) makes use of the positive definiteness of the conditional covariance matrix $H(S^{\prime}JZ_{t-1})$ (or $\sum_{m=1}^{M}\alpha_{m,t}\Omega_{m}$ in our notation) to conclude that the smallest eigenvalue of $H(S^{\prime}Jx)$, $x\in\mathbb{R}^{dp}$, is bounded away from zero in compact subsets of $\mathbb{R}^{dp}$. Saikkonen (2008) argues that thereby by Theorem 2.2(ii) of Cline and Phu (1998), the Markov chain $Z_{t}$ is an irreducible and aperiodic T-chain on $\mathbb{R}^{dp}$, which is then used to conclude further results.
Concerning our model, note that by Assumption B.1(III), the regime weights are either of the threshold or logistic form. In the threshold case, denoting a generic lag vector as $\boldsymbol{y}\in\mathbb{R}^{dp}$, the impact matrix $B_{y,t}(\boldsymbol{y})=\sum_{m=1}^{M}\alpha_{m,t}(\boldsymbol{y})B_{m}$ is exactly $B_{m}$ within each regime, and since each $B_{m}$ is assumed invertible, the conditional covariance matrix $B_{y,t}(\boldsymbol{y})B_{y,t}(\boldsymbol{y})$’ is positive definite. In the logistic two-regime case, Condition (IV) rules out singular convex combinations of $B_{1}$ and $B_{2}$, which likewise guarantees that $B_{y,t}(\boldsymbol{y})$ is invertible for all $\boldsymbol{y}$. [^13] Indeed, denoting $C\equiv B_{1}^{-1}B_{2}$ and $B(\alpha)=(1-\alpha)B_{1}+\alpha B_{2}=B_{1}((1-\alpha)I+\alpha C)$, the following conditions are equivalent: (1) $B(\alpha)$ is singular for some $\alpha^{*}\in(0,1)$ and (2) $C$ has a negative real eigenvalue. To prove this, observe first that $\det B(\alpha)=\det(B_{1})\det((1-\alpha)I+\alpha C)$. As $B_{1}$ is invertible, it follows that $B(\alpha)$ is singular iff $(1-\alpha)I+\alpha C$ is singular, i.e., iff there exists $v\neq 0$ with $((1-\alpha)I+\alpha C)v=0$, which is equivalent to $Cv=-\frac{1-\alpha}{\alpha}v$. For $\alpha\in(0,1)$, the scalar $\frac{1-\alpha}{\alpha}$ is real and negative. Hence, if $B(\alpha^{*})$ is singular for some $\alpha^{*}\in(0,1)$, then $C$ has a negative real eigenvalue $\tfrac{1-\alpha^{*}}{\alpha^{*}}$. Conversely, if $Cv=\lambda v$ with $\lambda<0$, choose $\alpha^{*}=\frac{1}{1-\lambda}\in(0,1)$. Then, $(1-\alpha^{*})+\alpha^{*}\lambda=0$, so $(1-\alpha^{*})I+\alpha^{*}C$ (and hence $B(\alpha^{*}))$ is singular, establishing the equivalence. Therefore, and because in the logistic case $B_{y,t}(\boldsymbol{y})$ is continuous in $\boldsymbol{y}$, the smallest eigenvalue of $B_{y,t}(\boldsymbol{y})B_{y,t}(\boldsymbol{y})$’ attains a strictly positive minimum on every compact subset of $\mathbb{R}^{dp}$. Thus, the conditional covariance matrix is uniformly bounded away from singularity on compact subsets as required. It follows from this, Equation (3.2), and the invertibility of $B_{y,t}$ that we can apply Theorem 2.2(ii) of Cline and Phu (1998) to conclude that the Markov chain $\boldsymbol{y}_{t}$ is an irreducible and aperiodic T-chain on $\mathbb{R}^{dp}$, similarly to the process $Z_{t-1}$ in Saikkonen (2008).
[…]
@@
$\sup\{||A||:A\in\cup_{j=1}^{N}\mathcal{A}^{j}\}<c,$ (C.37)
where $||\cdot||$ is Euclidean norm, $N$ is a positive integer, and $c$ a finite real number. By Jensen’s inequality and submultiplicativity of norms, we have
$E\left[\sup_{A\in\mathcal{A}^{j}}||AB_{y,t}(\boldsymbol{y})e_{t}||^{2}\middle|\boldsymbol{y}=x\right]\leq\sup_{A\in\mathcal{A}^{j}}||A||^{2}||B_{y,t}(x)||^{2}E||e_{t}||^{2},$ (C.38)
where the right side is bounded from above by some finite constant (by (C.37) and the finite second moment of $e_{t}$ given by Assumption 1). Thus, an appropriate version of Equation (A.5) in Saikkonen (2008) holds for our model. Since the rest of the proof of Theorem 1 in Saikkonen (2008) thereby applies to our model as well, this concludes the proof.∎
Appendix D Penalized nonlinear least squares estimation of STVAR models
@@
D.1 Nonlinear least squares estimation
@@
$Q(\alpha)=\sum_{t=1}^{T}u_{t}(\alpha)^{\prime}u_{t}(\alpha),$ (D.1)
@@
where $u_{t}(\alpha)=y_{t}-\sum_{m=1}^{M}\alpha_{m,t}\mu_{m,t}$ and the parameter $\alpha$ contains the weight function parameters. [^14] With threshold weights (A.1), $\alpha=(r_{1},...,r_{M-1})$, and with logistic weights (2.3) weights, $\alpha=(c,\gamma)$. If the transition weights are exogenous, we simply treat them as known constants. The NLS estimates of the parameters $\phi_{m},A_{m,1},...,A_{m,p}$, m=1,,Mm=1,...,M, corresponding to the weight function parameter $\alpha$ are obtained by minimizing $Q(\alpha)$ with respect to these parameters.
@@
If the transition weights $\alpha_{1,t},...,\alpha_{M,t}$ are not exogenous, also the parameter $\alpha$ they depend on needs to be estimated. Since the NLS estimates of the AR parameters are analytically obtained for given $\alpha$, the optimization can be performed with respect to $\alpha$ only. However, this optimization problem is multimodal, and thereby standard gradient based optimization methods that are not able escape from local maxima are insufficient. Since we only need approximate estimates to serve as starting values for Step 2 and 3 estimation in the three-step procedure described in Section 4.3, the following procedure is adopted.
@@
First, determine the values of the transition weights $\alpha_{m,t}$, m=1,,Mm=1,...,M, t=1,,Tt=1,...,T, that can be deemed reasonable for estimation, i.e., transition weights that allocate a sufficient enough contribution to each regime in the model. As an extreme example, transition weights with $\alpha_{1,t}=0$ for all t=1,,Tt=1,...,T are not reasonable, as Regime 1 would not be involved in the model at all. Thus, each regime should contribute to the model to an extend that is reasonable enough for estimation.
@@
One way to assess the contribution of each regime is by summing over the transition weights, i.e., calculating the sums $\dot{\alpha}_{m}\equiv\sum_{t=1}^{T}\alpha_{m,t}$, m=1,,Mm=1,...,M. In the special case of discrete transition weights such as the threshold weights (A.1), the sums $\dot{\alpha}_{m}$ express the numbers of observations from each regime. It can then be determined whether each regime contributes enough to the model by checking, for all m=1,,Mm=1,...,M, whether the sums $\dot{\alpha}_{m}$ are large enough compared to the number of parameters in that regime. Denoting the large enough sum as $\tilde{T}$, transition weights that satisfy $\dot{\alpha}_{m}\geq\tilde{T}$ for all m=1,,Mm=1,...,M are deemed reasonable enough. [^15] In our empirical application (Section 5), we assume that there should be at least three times as many observations from all the $d$ variables combined, implying that $\tilde{T}=3kd^{-1}$, where $k$ is the number of parameters in a regime. We acknowledge that this assumption is somewhat arbitrary, but it attempts to deter obvious cases of overfitting, while still allowing flexibility when the model is large compared to the number of observations.
@@
Second, specify a range of reasonable parameter values for each scalar parameter in the $(s\times 1)$ vector $\alpha$, and specify a grid of parameter values within that range. For instance, for a logistic STVAR model, the location parameter should, at least, lie between the lowest and greatest value in the observed series of the switching variable, whereas the scale parameter can range from a small positive number to some large positive number that implies almost discrete regime-switches. With $n$ grid points for each $s$ scalar parameters in $\alpha$, the number of different parameter vectors $\alpha$ with values in the grid points is $n^{s}$. Hence, we recommend selecting a sufficiently small number of grid points $n$, so that estimation can be performed in a reasonable computation time. Denote the $i$th vector of $\alpha$ containing parameter values in the grid points as $\alpha^{(i)}$, $i=1,...,n^{s}$. Then, obtain the subset of the grid points $\alpha^{(i)}$ that satisfy $\dot{\alpha}_{m}\geq\tilde{T}$ for all m=1,,Mm=1,...,M. Denote this subset as $\mathcal{W}$.
@@
Finally, calculate the NLS estimates of the AR parameters for all $\alpha^{(i)}\in\mathcal{W}$, and the estimates yielding the smallest sum of squares of residuals $Q(\alpha)$ are the approximate NLS estimates of the parameters $\phi_{m},A_{m,1},...,A_{m,p}$, m=1,,Mm=1,...,M, and $\alpha$. Notably, the NLS estimation does not necessarily produce estimates that satisfy the usual stability condition for all or any of the regimes.
D.2 Penalized nonlinear least squares estimation
@@
Penalized nonlinear least squares estimation of STVAR models works similarly to the NLS estimation, except that instead of the sum of squares function (D.1), the penalized sum of squares function is minimized. We define the penalized sum of squares function as
$PQ(\alpha)=Q(\alpha)-P(\alpha),$ (D.2)
@@
$P(\alpha)=\kappa\widehat{RSS}\sum_{m=1}^{M}\sum_{i=1}^{dp}\text{max}\{0,|\rho(\boldsymbol{A}_{m}(\boldsymbol{\theta}))_{i}|-(1-\eta)\}^{2},$ (D.3)
@@
$\widehat{RSS}=\underset{\alpha\in\mathcal{W}}{\min}Q(\alpha)$ (D.4)
@@
is the minimized residual sum of squares, $|\rho(\boldsymbol{A}_{m}(\boldsymbol{\theta}))_{i}|$ is the modulus of the $i$th eigenvalue of the companion form AR matrix of regime $m$, the tuning parameter $\eta\in(0,1)$ determines how close to the boundary of the stability region the penalization starts, and $\kappa>0$ determines the strength of the penalization. Whenever the companion form AR matrix of a regime has eigenvalues greater than $1-\eta$ in modulus, the penalization term (D.3) is greater than zero, and it increases in the modulus of these eigenvalues. The penalty term (D.3) is multiplied by $\widehat{RSS}$ to standardize the strength of the penalization with respect to the fit of the model.
@@
The penalized NLS estimates of $\phi_{m},A_{m,1},...,A_{m,p}$, m=1,,Mm=1,...,M, and $\alpha$, are obtained by first minimizing $Q(\alpha)$ for all $\alpha\in\mathcal{W}$ (where $\mathcal{W}$ is the set of grid point defined in Section D.1), then calculating the penalty term $P(\alpha)$ for all $\alpha\in\mathcal{W}$, and finally selecting the estimates that give the smallest minimized the penalized sum of squares (D.2). In our empirical analysis in Section 5, we use the tuning parameter values $\eta=0.05$ and $\kappa=0.2$, in line with the tuning parameter values used in the PML estimation (see Section 4.2).
Appendix E A Monte Carlo study